OpenStudy (anonymous):

\[Evaluate \int\limits x ^{3}\cos ^{5}x ^{4}sinx ^{4}dx\]

OpenStudy (lgbasallote):

try starting with let u = x^4 du = 4x^3 so the integral becomes \[\frac 14 \int \cos^5 (u) \sin (u )du\] does that help?

OpenStudy (anonymous):

yes. thanks. :)